OpenStudy (anonymous):

Simplify. your answer should contain only positive exponents 2u. u^2/3

OpenStudy (anonymous):

\[2\times u \times u ^{2/3}=2 \times u ^{1+2/3} = 2 \times u^{5/3}\]

OpenStudy (anonymous):

so which is the answeR?

OpenStudy (anonymous):

as per the question presented it's the most simplified one I can do.